The period of a helical spring is given by T=2pie, assume that e=50.0+-0.2(s.e). If time for 15 oscillations are as recorded below:24.2,24.2,24.7,24.5,24.2,24.2,24.2,24.4,24.5,24.6,24.7,24.5,24.6,24.5,24.3. Determine the value of g and the standard error in g


1
Expert's answer
2021-02-22T10:22:24-0500
"T=24.42\ s\\SE=\frac{s}{\sqrt{n}}=\frac{0.1897}{\sqrt{15}}=0.05\ s"